Show that (0,0) is a critical point of f(x, y) = x ^ 2 + kxy + y ^ 2 no matter what value the


constant khas. (Hint: Consider two cases: k = 0 and k * 0.)


1
Expert's answer
2021-11-01T20:03:07-0400

In the fisrt case, let k not equal to 0. So we get the following fx(x,y)=2x+ky=0 and fy(x,y)=kx+2y=0 therefore, fx=0 implies y=-(2x)/(k) which also implies kx+2(-(2x)/(k))=kx-(4x)/(k)=x(k-(4)/((k)))=0

This implies x=0Vk=+-2 which also implies 

y=0Vy=+-x

In both cases you get the point (0,0) which is a critical point.

If we let k=0,

We get f(x,y)=x^(2)+y^(2) which implies fx(x,y)=2x=0

fy(x,y)=2y=0 implies that x=0 and y=0

In both cases, the critical point is (0,0)